# Segmenting Action-Value Functions Over Time-Scales in SARSA using TD($\Delta$)

### v1

@@ Abstract
In numerous episodic reinforcement learning (RL) settings, SARSA-based methodologies are employed to enhance policies aimed at maximizing returns over long horizons. Conventional SARSA algorithms, however, have difficulties in balancing bias and variation due to the reliance on a singular, fixed discount factor. This study expands the temporal difference decomposition approach, TD($\triangle$), to the SARSA algorithm. SARSA, a widely utilised on-policy RL method, enhances action-value functions via temporal difference updates. TD($\triangle$) facilitates learning over several time-scales by breaking the action-value function into components associated with distinct discount factors. This decomposition improves learning efficiency and stability, particularly in problems necessitating long-horizon optimization. We illustrate that our methodology mitigates bias in SARSA’s updates while facilitating accelerated convergence in contexts characterized by dense rewards. Experimental findings across many benchmark tasks indicate that the proposed SARSA($\triangle$) surpasses conventional TD learning methods in both tabular and deep RL contexts.
@@ 1 Introduction
In reinforcement learning (RL), the objective is to enhance an agent’s actions to maximize cumulative rewards through interaction with an environment. Temporal difference (TD) learning methods, like Q-learning and SARSA, have demonstrated efficacy across several tasks by allowing agents to estimate action-value functions that forecast anticipated future rewards. Traditionally, these strategies utilise a discount factor $10\leq\gamma<$ 1, with values approaching $\gamma=0$ to emphasise short-term rewards over long-term ones, therefore diminishing the planning horizon and enhancing learning stability and efficiency[1]. Discount factors $\gamma<1$ frequently yield enhanced results during the early phases of learning, as demonstrated by Prokhorov and Wunsch[2]. In specific contexts, such as long-horizon tasks[3, 4] where long-term planning is essential—like navigation or decision-making tasks with deferred rewards—this method can create bias, constraining the agent’s capacity to acquire optimal long-horizon policies.
@@
For instance: SARSA[1] (State-Action-Reward-State-Action) is an on-policy RL algorithm that modifies an action-value function Q(s,a) according to the state-action transitions encountered by the agent. The conventional SARSA algorithm performs well in contexts with shorter time horizons; however, in long-horizon tasks, the singular discount factor $\gamma$ employed in SARSA results in a compromise between bias and variance. Small values of $\gamma$ yield myopic policies, whereas larger values induce substantial variance in the estimated action-value functions, rendering the learning process unstable and inefficient. In response, other solutions have been suggested, including[4, 2, 5, 6, 7, 8].
@@
Recent research in TD learning has investigated the segmentation of value functions over many time-scales, enabling agents to concurrently enhance value estimates at numerous horizons[8]. The TD($\Delta$) architecture enhances scalability and performance in situations with long-term dependencies by decomposing value functions into components that encompass various discount factors, referred to as TD($\Delta$). The TD($\Delta$) framework’s adaptability demonstrates significant potential in mitigating the bias-variance trade-offs inherent in TD learning.
@@
We suggest SARSA($\Delta$), an extension of the TD($\Delta$) method applied to the SARSA algorithm, building on this concept. TD($\Delta$) emphasises the actor-critic paradigm and conventional TD learning, whereas SARSA($\Delta$) disaggregates the state-action value function into several partial estimators across various discount factors, termed delta estimators. These estimators approximate the difference $W_{z}(s,a)=Q_{\gamma_{z}}(s,a)\mathit{-}Q_{\gamma_{z-1}}(s,a)$ between action-value functions, facilitating efficient learning in scenarios where actions affect long-horizon outcomes. SARSA($\Delta$) optimises a sequence of delta estimators, each associated with a distinct discount factor, similar to TD($\Delta$). The segregation enables the autonomous learning of each component, with lower discount factors achieving faster convergence, while bigger discount factors enhance this foundation for improved long-horizon planning.
@@
This study introduces a multi-step SARSA($\Delta$) update rule that disaggregates state-action value functions, facilitating accelerated convergence and enhanced policy efficacy, especially in intricate environments characterised by dense rewards. Furthermore, we illustrate that SARSA($\Delta$) may seamlessly adapt to any multi-step reinforcement learning techniques, including n-step SARSA and eligibility traces. Through theoretical analysis and empirical findings, we emphasise the benefits of SARSA($\Delta$) in diverse tasks, including a basic ring MDP utilised by Kearns & Singh[9], demonstrating its superiority over conventional SARSA in scenarios necessitating a balance between short-term action consequences and long-term reward maximisation. SARSA($\Delta$) possesses the capability to adapt discount factors across varying time scales, serving as an effective instrument in environments where time scale is critical.
@@
The primary contributions of this paper are as follows: (i) We expand TD($\Delta$) to SARSA($\Delta$) by partitioning the action-value function Q(s,a) into delta components over several time scales. (ii) The suggested SARSA($\Delta$) algorithm enhances learning efficiency and stability by independently learning components for various discount factors and integrating them to represent the complete action-value function. (iii) Both theoretical and empirical evaluations have shown that this multi-scale breakdown has a number of benefits, particularly in environments where there are long-horizon rewards.
The subsequent sections of the paper are structured as follows: Section 2 contains a discussion of related works. Section 3 presents the necessary background information. Sections 4 and 5 illustrate the principles of SARSA($\Delta$) and theoretical analysis, respectively. Section 6 presents a comprehensive overview of the experiments conducted. Ultimately, we provide a conclusion in section 7.
@@ 2 Related Work
In reinforcement learning (RL), optimising for long-horizon rewards is tough due to the complexity of learning with an undiscounted return. Temporal discounting is frequently employed to streamline this process; nevertheless, it may induce bias. Researchers[8] have investigated techniques to separate value functions across various time scales by analysing disparities among value functions with lower discount factors, hence enhancing scalability and performance. The study introduces $\gamma$-nets[10], a technique for generalising value function estimate across timescales in reinforcement learning. The authors of the paper[11] present an approach for integrating hyperbolic discounting into RL. Traditional RL employs exponential discounting, which does not align with the hyperbolic discounting observed in human and animal behaviors. The authors present an agent designed to approximate hyperbolic discount functions by utilizing temporal-difference learning methods. Furthermore, the study reveals that learning value functions across multiple time horizons improves performance, especially when implemented as an auxiliary task in value-based reinforcement learning algorithms such as Rainbow. This method demonstrates potential in contexts characterized by uncertainty and risk. Expected SARSA[7], a variant of SARSA, utilizes information regarding the stochastic nature of the behavior policy to execute updates with reduced variance, which speeds up learning. This study[12] presents temporal-difference (TD) learning methods, which underpin numerous RL algorithms, notably SARSA, by forecasting future outcomes based on the discrepancies between consecutive forecasts rather than the ultimate result. The research demonstrates the convergence of TD techniques under specific situations and emphasises their benefits in predictive tasks, particularly when incremental learning is crucial. This study[13] presents theoretical convergence results for temporal-difference learning methods utilising function approximation. Extending this analysis to SARSA($\Delta$) enables us to assert comparable convergence properties, provided that the delta components are updated in accordance with the SARSA update rule. It is noteworthy that none of the previously mentioned works employ short-term estimates for training long-term action-value functions. Our method can also be utilised as a generalised action-value function, and its capacity to query smaller timescales represents an additional advantage. Performance improvements primarily arise from the separation of action-value functions across varying time scales through the use of TD($\Delta$) in SARSA.
@@
Lastly, another extensive category of study pertinent to our research indirectly encompasses hierarchical RL and the bias-variance tradeoff. These studies[14, 15, 16, 17, 18, 19, 20] present the concept of deconstructing value functions into subtasks by hierarchical RL, analogous to the manner in which TD($\Delta$) disaggregates value functions across temporal scales. This substantiates the assertion that value function decomposition can enhance learning efficacy and stability. This paper[9] functions as a theoretical foundation for comprehending the bias-variance tradeoff that arises when decomposing Q-values across various time scales, and it also establishes constraints on the bias and variance of TD updates. This study[21] presents empirical findings demonstrating the efficacy of deep RL techniques utilizing value function approximation in Atari-like environments. It also endorses the necessity for techniques that diminish variance in long-horizon tasks.
@@ 3 Background and notation
Examine a completely observable Markov Decision Process (MDP) [22], characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ signifies the action space, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\rightarrow[0,1]$ represents the transition probabilities that associate state-action pairings with distributions across subsequent states, whereas $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function. At each timestep $t$, the agent occupies state $s_{t}$, chooses an action $a_{t}$, obtains a reward $r_{t}=r(s_{t},a_{t})$, and transitions to the subsequent state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$.
@@
In a conventional Markov Decision Process (MDP) framework, an agent seeks to maximize the discounted return defined as $Q_{\gamma}^{\pi}(s,a)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t+1}\mid s_{t}=s,a_{t}=a\right]$, where $\gamma$ denotes the discount factor and $\pi:\mathcal{S}\rightarrow\mathcal{A}\rightarrow[0,1]$ represents the policy adhered to by the agent. The action-value function $Q_{\gamma}^{\pi}(s,a)$ is determined as the fixed point of the Bellman operator $\mathcal{T}Q=r^{\pi}+\gamma\mathcal{P}^{\pi}Q$, where $r^{\pi}$ denotes the expected immediate reward and $\mathcal{P}^{\pi}$ represents the transition probability operator associated with the policy $\pi$. For convenience, we remove the superscript $\pi$ for the rest of the paper.
@@
Using temporal difference (TD)[23] learning, the action-value estimate $\hat{Q}_{\gamma}$ can approximate the true action-value function $Q_{\gamma}$. The one-step TD error $\delta^{\gamma}_{t}=r_{t+1}+\gamma\hat{Q}_{\gamma}(s_{t+1},a_{t+1})\mathit{-}\hat{Q}_{\gamma}(s_{t},a_{t})$ is used to update the action-value function given a transition $(s_{t},a_{t},r_{t},s_{t+1})$.
SARSA is an on-policy RL approach that acquires the action-value function Q(s,a) for a specified policy $\pi$. The action-value function denotes the anticipated cumulative reward originating from state s, executing action a, and following policy $\pi$. Subsequently, the SARSA update rule is articulated as follows:
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left[r_{t}+\gamma Q(s_{t+1},a_{t+1})-Q(s_{t},a_{t})\right]$ (1)
@@
where $\alpha$ denotes the learning rate, $\gamma$ represents the discount factor, and $r_{t}$ signifies the immediate reward. Long-horizon tasks present challenges for SARSA since the choice of $\gamma$ can compromise learning efficiency with appropriate long-term reward maximizing.
@@
On the other hand, for a complete trajectory, we can employ the discounted sum of one-step TD errors, generally known as the $\lambda$-return[23] or, equivalently, the Generalized Advantage Estimator (GAE)[24]. The GAE enhances advantage estimates by reconciling the trade-off between variance and bias via parameters $\lambda$ and $\gamma$. The equation for the Generalized Advantage Estimator is generally expressed as:
$A(s_{t},a_{t})=\sum_{k=0}^{\infty}(\lambda\gamma)^{k}\delta_{t+k}^{\gamma}$ (2)
Where $\delta_{t+k}$ is the TD error at time t, computed as follows: $\delta_{t+k}=r_{t+k}+\gamma Q(s_{t+k+1},a_{t+k+1})-Q(s_{t+k},a_{t+k})$
@@
Loss function for Q-Value estimation utilizing GAE. The loss function $\mathcal{L}(\theta)$ for approximating the Q-value function is defined as the mean squared error between the current Q-value estimate $Q(s_{t},a_{t};\theta)$ and the target Q-value adjusted by the advantage estimator. Thus, it is possible to concisely write the loss function using $A(s_{t},a_{t})$ and $Q(s_{t},a_{t})$ as follows:
$\mathcal{L}(\theta)=\mathbb{E}\bigg{[}\bigg{(}Q(s_{t},a_{t};\theta)-\bigg{(}Q(s_{t},a_{t})+A(s_{t},a_{t})\bigg{)}\bigg{)}^{2}\bigg{]}$ (3)
@@
Despite SARSA being on-policy and lacking an explicit policy update, policy selection in SARSA can still be influenced by Eq.4. For actor-critic architectures[25, 26, 27], the action-value function is updated according to Eq.3, and a stochastic parameterized policy (actor, $\pi_{\omega}(a|s)$) is learned from this value estimator through the advantage function, with the loss being.
$\mathcal{L}(\omega)=\mathbb{E}\bigg{[}-log\pi(a|s;\omega)A(s,a)\bigg{]}$ (4)
@@
In on-policy methods like SARSA, the likelihood ratio $\rho$ between the current policy and a prior policy is typically not employed. SARSA evaluates actions in accordance with the current policy, concurrently updating the action-value function as actions are performed. Thus, Proximal Policy Optimization (PPO)[28] does not implement a clipping objective function in SARSA as used in this study[8]. Eq.4 is utilized as the objective function.

### v2

# Segmenting Action-Value Functions Over Time-Scales in SARSA via TD($\Delta$)

@@ Abstract
In numerous episodic reinforcement learning (RL) settings, SARSA-based methodologies are employed to enhance policies aimed at maximizing returns over long horizons. Conventional SARSA algorithms, however, have difficulties in balancing bias and variation due to the reliance on a singular, fixed discount factor. This study expands the temporal difference decomposition approach, TD($\Delta$), to the SARSA algorithm, which we designate as SARSA($\Delta$). SARSA, a widely utilised on-policy RL method, enhances action-value functions via temporal difference updates. TD($\Delta$) facilitates learning over several time-scales by breaking the action-value function into components associated with distinct discount factors. This decomposition improves learning efficiency and stability, particularly in problems necessitating long-horizon optimization. We illustrate that our methodology mitigates bias in SARSA’s updates while facilitating accelerated convergence in both deterministic and stochastic environments. Experimental findings across many benchmark tasks indicate that the proposed SARSA($\Delta$) surpasses conventional TD learning methods in both tabular and deep RL environments.
@@ 1 Introduction
In reinforcement learning (RL), the objective is to enhance an agent’s actions to maximize cumulative rewards through interaction with an environment. Temporal difference (TD) learning methods, like Q-learning and SARSA, have demonstrated efficacy across several tasks by allowing agents to estimate action-value functions that forecast anticipated future rewards. Traditionally, these strategies utilise a discount factor $10\leq\eta<$ 1, with values approaching $\eta=0$ to emphasise short-term rewards over long-term ones, therefore diminishing the planning horizon and enhancing learning stability and efficiency[1]. Discount factors $\eta<1$ frequently yield enhanced results during the early phases of learning, as demonstrated by Prokhorov and Wunsch[2]. In specific contexts, such as long-horizon tasks[3, 4] where long-term planning is essential—like navigation or decision-making tasks with deferred rewards—this method can create bias, constraining the agent’s capacity to acquire optimal long-horizon policies.
@@
For instance: SARSA[1] (State-Action-Reward-State-Action) is an on-policy RL algorithm that modifies an action-value function Q(s,a) according to the state-action transitions encountered by the agent. The conventional SARSA algorithm performs well in contexts with shorter time horizons; however, in long-horizon tasks, the singular discount factor $\eta$ employed in SARSA results in a compromise between bias and variance. Small values of $\eta$ yield myopic policies, whereas larger values induce substantial variance in the estimated action-value functions, rendering the learning process unstable and inefficient. In response, other solutions have been suggested, including[4, 2, 5, 6, 7, 8].
@@
Recent research in TD learning has investigated the segmentation of value functions across multiple time scales, enabling agents to enhance value estimates concurrently at different horizons[8]. The TD($\Delta$) architecture improves scalability and performance in scenarios involving long-term dependencies by breaking down value functions into components that incorporate different discount factors, known as TD($\Delta$). The adaptability of the TD($\Delta$) framework shows considerable promise in addressing the bias-variance trade-offs that are characteristic of TD learning.
@@
We propose SARSA($\Delta$), an enhancement of the TD($\Delta$) approach integrated into the SARSA algorithm, based on this principle. TD($\Delta$) emphasizes the actor-critic framework and conventional TD learning, whereas SARSA($\Delta$) divides the state-action value function into multiple partial estimators corresponding to various discount factors, known as delta estimators. These estimators approximate the difference $D_{m}(s,a)=Q_{\eta_{m}}(s,a)-Q_{\eta_{m-1}}(s,a)$ between action-value functions, thereby improving learning efficiency in scenarios where actions affect long-horizon outcomes. Similar to TD($\Delta$), SARSA($\Delta$) seeks to optimize a set of delta estimators, where each estimator is associated with a distinct discount factor. With smaller discount factors, convergence happens faster because to segmentation, while greater discount factors strengthen this foundation for better long-horizon planning.
@@
This study introduces a multi-step SARSA($\Delta$) update rule that disaggregates action-value functions, facilitating accelerated convergence and enhanced policy efficacy, especially in intricate deterministic and stochastic environments. Furthermore, we illustrate that SARSA($\Delta$) may seamlessly adapt to any multi-step reinforcement learning techniques, including n-step SARSA and eligibility traces. Through theoretical analysis and empirical findings, we emphasise the benefits of SARSA($\Delta$) in diverse tasks, including a basic ring MDP utilised by Kearns & Singh[9], demonstrating its superiority over conventional SARSA in scenarios necessitating a balance between short-term action consequences and long-term reward maximisation. SARSA($\Delta$) possesses the capability to adapt discount factors across varying time scales, serving as an effective instrument in environments where time scale is critical.
@@
The primary contributions of this paper are as follows: (i) We expand TD($\Delta$) to SARSA($\Delta$) by partitioning the action-value function Q(s,a) into delta components over several time-scales. (ii) The suggested SARSA($\Delta$) algorithm enhances learning efficiency and stability by independently learning components for various discount factors and integrating them to represent the complete action-value function. (iii) Both theoretical and empirical evaluations have shown that this multi-scale breakdown has a number of benefits, particularly in environments where there are long-horizon rewards.
The ensuing portions of the paper are organized as follows: Section 2 presents a discourse on pertinent literature. Section 3 provides the essential background information. The principles of SARSA($\Delta$) and theoretical analysis are shown in Sections 4 and 5, respectively. Section 6 provides a detailed summary of the experiments performed. Ultimately, we present a conclusion in section 7.
@@ 2 Related Work
In reinforcement learning (RL), optimising for long-horizon rewards is tough due to the complexity of learning with an undiscounted return. Temporal discounting is frequently employed to streamline this process; nevertheless, it may induce bias. Researchers[8] have investigated techniques to separate value functions across various time scales by analysing disparities among value functions with lower discount factors, hence enhancing scalability and performance. The study introduces $\gamma$-nets[10], a technique for generalising value function estimate across timescales in RL. The authors of the paper[11] present an approach for integrating hyperbolic discounting into RL. Traditional RL employs exponential discounting, which does not align with the hyperbolic discounting observed in human and animal behaviors. The authors present an agent designed to approximate hyperbolic discount functions by utilizing temporal-difference learning methods. Furthermore, the study reveals that learning value functions across multiple time horizons improves performance, especially when implemented as an auxiliary task in value-based RL algorithms such as Rainbow. This method demonstrates potential in contexts characterized by uncertainty and risk. Expected SARSA[7], a variant of SARSA, utilizes information regarding the stochastic nature of the behavior policy to execute updates with reduced variance, which speeds up learning. This study[12] presents temporal-difference (TD) learning methods, which underpin numerous RL algorithms, notably SARSA, by forecasting future outcomes based on the discrepancies between consecutive forecasts rather than the ultimate result. The research demonstrates the convergence of TD techniques under specific situations and emphasises their benefits in predictive tasks, particularly when incremental learning is crucial. This study[13] presents theoretical convergence results for temporal-difference learning methods utilising function approximation. Extending this analysis to SARSA($\Delta$) enables us to assert comparable convergence properties, provided that the delta components are updated in accordance with the SARSA update rule. It is noteworthy that none of the previously mentioned works employ short-term estimates for training long-term action-value functions. Our method can also be utilised as a generalised action-value function, and its capacity to query smaller time-scales represents an additional advantage. Performance improvements primarily arise from the separation of action-value functions across varying time scales through the use of TD($\Delta$) in SARSA.
@@
Lastly, another extensive category of study pertinent to our research indirectly encompasses hierarchical RL and the bias-variance trade-off. These studies[14, 15, 16, 17, 18, 19, 20] present the concept of deconstructing value functions into subtasks by hierarchical RL, analogous to the manner in which TD($\Delta$) disaggregates value functions across temporal scales. This substantiates the assertion that value function decomposition can enhance learning efficacy and stability. This paper[9] functions as a theoretical foundation for comprehending the bias-variance trade-off that arises when decomposing Q-values across various time-scales, and it also establishes constraints on the bias and variance of TD updates. This study[21] presents empirical findings demonstrating the efficacy of deep RL techniques utilizing value function approximation in Atari-like environments. It also endorses the necessity for techniques that diminish variance in long-horizon tasks.
@@ 3 Background and notation
Examine a completely observable Markov Decision Process (MDP) [22], characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ signifies the action space, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\rightarrow[0,1]$ represents the transition probabilities that associate state-action pairings with distributions across subsequent states, whereas $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function. At each timestep $n$, the agent occupies state $s_{n}$, chooses an action $a_{n}$, obtains a reward $r_{n}=r(s_{n},a_{n})$, and transitions to the subsequent state $s_{n+1}\sim\mathcal{P}(\cdot|s_{n},a_{n})$.
@@
In a conventional Markov Decision Process (MDP) framework, an agent seeks to maximize the discounted return defined as $Q_{\eta}^{\pi}(s,a)=\mathbb{E}\left[\sum_{n=0}^{\infty}\eta^{n}r_{n+1}\mid s_{n}=s,a_{n}=a\right]$, where $\eta$ denotes the discount factor and $\pi:\mathcal{S}\rightarrow\mathcal{A}\rightarrow[0,1]$ represents the policy adhered to by the agent. The action-value function $Q_{\eta}^{\pi}(s,a)$ is determined as the fixed point of the Bellman operator $\mathcal{T}Q=r^{\pi}+\eta\mathcal{P}^{\pi}Q$, where $r^{\pi}$ denotes the expected immediate reward and $\mathcal{P}^{\pi}$ represents the transition probability operator associated with the policy $\pi$. For convenience, we remove the superscript $\pi$ for the rest of the paper.
@@
Using temporal difference (TD)[23] learning, the action-value estimate $\hat{Q}_{\eta}$ can approximate the true action-value function $Q_{\eta}$. The one-step TD error $\delta^{\eta}_{n}=r_{n+1}+\eta\hat{Q}_{\eta}(s_{n+1},a_{n+1})\mathit{-}\hat{Q}_{\eta}(s_{n},a_{n})$ is used to update the action-value function given a transition $(s_{n},a_{n},r_{n},s_{n+1})$.
An on-policy RL technique called SARSA learns the action-value function Q(s,a) for a given policy $\pi$. The action-value function represents the expected total reward that comes from state s, taking action a, and following policy $\pi$. Then the SARSA update rule is clarified as follows:
$Q(s_{n},a_{n})\leftarrow Q(s_{n},a_{n})+\alpha\left[r_{n}+\eta Q(s_{n+1},a_{n+1})-Q(s_{n},a_{n})\right]$ (1)
@@
Here, $\alpha$ is the learning rate, $\eta$ is the discount factor, and $r_{n}$ is the immediate reward. Long-horizon tasks present challenges for SARSA since the choice of $\eta$ can compromise learning efficiency with appropriate long-term reward maximizing.
@@
On the other hand, for a complete trajectory, we can employ the discounted sum of one-step TD errors, generally known as the $\lambda$-return[23] or, equivalently, the Generalized Advantage Estimator (GAE)[24]. The GAE enhances advantage estimates by reconciling the trade-off between variance and bias via parameters $\lambda$ and $\eta$. The equation for the Generalized Advantage Estimator is generally expressed as:
$A(s_{n},a_{n})=\sum_{k=0}^{\infty}(\lambda\eta)^{k}\delta_{t+k}^{\eta}$ (2)
Where $\delta_{n+k}$ is the TD error at time n, computed as follows: $\delta_{n+k}=r_{n+k}+\eta Q(s_{n+k+1},a_{n+k+1})-Q(s_{n+k},a_{n+k})$
@@
Loss function for Q-Value estimation utilizing GAE. The loss function $\mathcal{L}(\theta)$ for approximating the Q-value function is defined as the mean squared error between the current Q-value estimate $Q(s_{n},a_{n};\theta)$ and the target Q-value adjusted by the advantage estimator. Thus, it is possible to concisely write the loss function using $A(s_{n},a_{n})$ and $Q(s_{n},a_{n})$ as follows:
$\mathcal{L}(\theta)=\mathbb{E}\bigg{[}\bigg{(}Q(s_{n},a_{n};\theta)-\bigg{(}Q(s_{n},a_{n})+A(s_{n},a_{n})\bigg{)}\bigg{)}^{2}\bigg{]}$ (3)
@@
Despite SARSA being on-policy and lacking an explicit policy update, policy selection in SARSA can still be influenced by Eq.4. For actor-critic architectures[25, 26, 27], the action-value function is updated according to Eq.3, and a stochastic parameterized policy (actor, $\pi_{\nu}(a|s)$) is learned from this value estimator through the advantage function, with the loss being.
$\mathcal{L}(\nu)=\mathbb{E}\bigg{[}-log\pi(a|s;\nu)A(s,a)\bigg{]}$ (4)
Proximal Policy Optimization (PPO) [28], an enhancement of actor-critic methodologies, confines policy updates to a designated optimization area, referred to as a trust region, utilizing a clipping objective that evaluates the current parameters, $\nu$, against the prior parameters, $\nu_{old}$:
$\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A(s,a),\psi(\nu)A(s,a)\bigg{)}\bigg{]}$ (5)
@@
In on-policy methods like SARSA, the likelihood ratio $\rho(\nu)$ between the current policy and a prior policy is typically employed. SARSA evaluates actions in accordance with the current policy, concurrently updating the action-value function as actions are performed. However, when utilized within an actor-critic framework, the SARSA-learning agent may adopt a similar policy ratio:
$\rho(\nu)=\frac{Q(s,a;\nu)}{Q(s,a;\nu_{old})}$ (6)
where $\psi(\nu)=clip(\rho,1-\epsilon,1+\epsilon)$ represents the clipped likelihood ratio, and $\epsilon<1$ is a negligible parameter employed to constrain the update. The loss function for the policy update can be articulated as follows:
$\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A(s,a),clip(\rho(\nu),1-\epsilon,1+\epsilon)A(s,a)\bigg{)}\bigg{]}$ (7)